# Explaining Time Series Classifiers with PHAR: Rule Extraction and Fusion from Post-hoc Attributions

### v3

@@ Abstract
Explaining machine learning (ML) models for time series (TS) classification remains challenging due to the difficulty of interpreting raw time series and the high dimensionality of the input space. We introduce PHARPost-hoc Attribution Rules—a unified framework that transforms numeric feature attributions from post-hoc, instance-wise explainers (e.g., LIME, SHAP) into structured, human-readable rules. These rules define human-readable intervals that indicate where and when decision-relevant segments occur and can enhance model transparency by localizing threshold-based conditions on the raw series. PHAR performs comparably to native rule-based methods, such as Anchor, while scaling more efficiently to long TS sequences and achieving broader instance coverage. A dedicated rule fusion step consolidates rule sets using strategies like weighted selection and lasso-based refinement, balancing key quality metrics: coverage, confidence, and simplicity. This fusion ensures each instance receives a concise and unambiguous rule, improving both explanation fidelity and consistency. We further introduce visualization techniques to illustrate specificity-generalization trade-offs in the derived rules. PHAR resolves conflicting and overlapping explanations—a common effect of the Rashomon phenomenon—into coherent, domain-adaptable insights. Comprehensive experiments on UCR/UEA Time Series Classification Archive demonstrate that PHAR may improve interpretability, decision transparency, and practical applicability for TS classification tasks by providing concise, human-readable rules aligned with model predictions.
@@ 1 Introduction
Explaining time series classifiers presents unique challenges compared to models for images or tabular data. Time series often lack obvious semantic features. Without domain expertise, even simple attribution outputs can be difficult to interpret, which complicates the use of standard Explainable AI (XAI) tools [37]. Temporal dependencies, phase shifts, and high dimensionality further worsens explanations. When non-specialized methods are applied, results can become unstable or overly generic [12]. These issues reveal a key gap in current XAI research. Unlike non-temporal models built on intuitive, semantically rich features, time series classifiers demand explanations that both capture temporal dynamics and deliver meaningful, context-aware insights tailored to sequential data. Moreover, deep learning (DL) is widely used for time series (TS) classification, particularly in medicine and industry, because of its ability to model complex temporal patterns for tasks such as disease diagnosis and predictive maintenance. Despite strong predictive performance, these models remain hard to interpret in practice.
@@
Aiming to address XAI challenges for state-of-the-art models such as DL, common approaches include numeric attribution methods like Permutation Importance [3], SHAP [17], and LIME [28], as well as gradient-based techniques like Saliency Maps [33] and Layer-wise Relevance Propagation (LRP) [5], all of which quantify feature contributions and trace relevance through the model’s computations. These approaches facilitate feature ranking, sensitivity analysis, and model debugging by producing developer-oriented outputs like importance scores or heatmaps. However, studies show that such numeric explanations often highlight broad regions of the input without delivering clear, context-specific insights [4, 15]. In sequential data, this tendency gives rise to "ambiguous rules", where multiple time steps or variables appear equally influential and cannot be easily distinguished. Consequently, these explanations are difficult for non-experts to interpret [14, 27], and the challenge is amplified in TS classifiers, where temporal dependencies and the high dimensionality of multivariate sequences further obscure meaningful patterns [30, 37].
@@
We preserve the strengths of numeric-based attribution methods: precise feature ranking, post-hoc interpretability, and local fidelity, while extending them to the time series classification domain. As a first step toward addressing these challenges, rule-based approaches provide a compact, human-readable alternative to numeric attributions. We employ semi-factual reasoning to identify minimal temporal segments—interpretable intervals where perturbing feature values would not change the model’s prediction—thus generating concise, human-readable rules. Anchor rule-based explanations have demonstrated improved interpretability by extracting succinct "IF–THEN" conditions that align closely with model behavior [29]. However, such methods are seldom applied to TS data. Furthermore, rule-based methods become computationally prohibitive as sequence length increases [1]. Exhaustive search for stable "IF–THEN" conditions could be computationally infeasible, leading to a combinatorial explosion of candidate rules and temporal instability. Beyond computational infeasibility, Cognitive Load Theory, summarized by [34], emphasizes reducing extraneous demands on human working memory. Analogously, simpler and unambiguous decision rules referencing a limited number of time steps minimize cognitive effort for experts, supporting interpretability and visualization by exposing time-localized, threshold-based conditions that can be inspected on the original signal.
Our motivation stems from the persistent difficulty of interpreting high-dimensional time series models with post-hoc methods, both attribution scores and rule-based. Moreover, existing work typically assesses single explainer separately, neglecting the advantages of fusing multiple sources and systematically evaluating their quality in terms of coverage, confidence, and simplicity. Finally, the opportunity to enhance transparency by overlaying explanations on time series plots has been largely overlooked. These gaps motivate our framework, which transforms instance-level attributions into concise, human-readable rules. Rather than relying on global XAI methods based on clustering or surrogate models [21], our approach leverages numeric attributions such as SHAP [17] and LIME [28]. We adopt Anchor [29] as our baseline for comparison. Our framework unifies insights from diverse explainers, measures rule quality, and visualizes explanations directly on time series to facilitate transparent decision-making. Our framework primarily targets TS but can extend to structured data like tabular datasets.
@@
PHAR aims to improve trust and transparency in TS classification with deep learning (DL) through three key contributions: (1) a pipeline that converts numeric post-hoc attributions into expert-level rule representations for TS; (2) a fusion mechanism that unifies rules from different explainers for better coherence and quality; (3) visualization of rules, as outlined in Figure 1. Each input time series is represented as a $T\times C$ matrix—$T$ time steps $t_{1},\dots,t_{T}$ by $C$ channels $c_{1},\dots,c_{C}$—and is processed by a deep learning classifier. Post-hoc explainers such as SHAP and LIME compute numerical attributions. In the first stage, these outputs are converted into interval rules for each instance. Rule-based explanations such as Anchor can be used directly. In the second stage, rule sets derived from different explainers are fused using methods such as union, weighted selection, and lasso-based refinement to yield a single, coherent rule per instance. Finally, each rule is visualized as vertical interval markers overlaid on the original time-series plot, enabling domain experts to inspect and validate the model’s decision logic. Crucially, the visualization conveys information beyond the textual rule: it shows rule–signal alignment in time, overlap across channels and explainers, which supports faster error checking and expert validation.
The remainder of this paper is structured as follows: Section 2 reviews related work and explains our motivation. Section 3 details our framework, covering numeric-to-rule transformation (Section 3.3) and fusion explainers (Section 3.4). Section 4 evaluates coverage, confidence, and other metrics. Section 5 summarizes findings and future directions. The Appendix A provides supplementary tables, figures, and extended results, while "Supplementary Materials" (Section B) contains statical results for Critical Difference Diagrams.
@@ 2 Related work
In this section, we review the state of the art in time series explainability that informs our PHAR framework. TS explanation methods often use surrogate models like decision trees for rule-based intervals. LIME [28], SHAP [17], and saliency maps offer numerical insights but fail to capture temporal dependencies in TS. DTW-based clustering identifies prototypical patterns, providing structural insights. Studies explore fuzzy rules, decision tree splits, and statistical impact rules, emphasizing interpretability. Some methods balance rule complexity, local fidelity, and temporal coherence, while others focus on local or global views, leaving gaps in integrating both effectively. We first examine rule-based explanation methods tailored to sequential data, and we then survey strategies for fusing multiple explainers.
@@
Explaining TS classifiers is challenging, especially for complex models.  [21] proposed a post-hoc method using Parameterized Event Primitives (PEPs) to explain TS models. Their approach detects key events (trends, extrema), clusters them, and derives rules via decision trees. These rules capture sequential patterns, supporting fidelity to model predictions and human interpretability, while acknowledging that such rules are proxies of model behavior rather than causal explanations. This method extracts meaningful patterns while maintaining temporal dependencies and lays the groundwork for hybrid explanations that address temporal dependencies.
@@
[32] propose a TS explanation framework combining saliency maps, instance-based examples, and rule-based methods. Saliency maps highlight key regions, instance-based methods provide prototypes and counterfactuals, and rule-based methods define logical conditions on subsequences. Integrating these methods bridges local and global explanations. This hybrid approach can improve interpretability and temporal coherence of the resulting explanations by constraining rules to compact time intervals and simple inequalities on the raw series.
@@
Decision rules enhance interpretability in TS forecasting, especially in finance.  [4] review decision trees, fuzzy trees, and Mamdani-type rules.  [31] introduced FDT-FTS, combining fuzzy logic and decision trees for multivariate TS forecasting. Mamdani-type rules use fuzzy logic and linguistic variables for human-readable explanations.  [40] integrated them into the Neural Fuzzy Hammerstein-Wiener network. These methods balance predictive accuracy with human-readable simplicity of rules (few conditions, short intervals), supporting TS analysis. Next, we discuss decision rules for progressive decision-making.
@@
The Belief–Rule–Based (BRB) framework [23] overcomes key weaknesses of post-hoc methods such as SHAP [17], LIME [28], and LINDA-BN [18] by structuring explanations into human-readable rules. In the Knowledge Graph Module (KGM), expert input is formalized into a causal graph that captures both temporal order and inter-feature dependencies, addressing SHAP’s and LIME’s tendency to treat each timestep or feature independently and thus lose important sequence information. The Knowledge Propagation Module (KPM) then propagates belief distributions through this graph to extrapolate feature trajectories, inherently managing missing data and uncertainty without relying on expensive local sampling that can render SHAP and LIME at times unstable. Finally, the Feature Evaluation Module (FEM) synthesizes these propagated beliefs into concise "IF–THEN" rules that highlight the most influential intervals, removing the need to interpret raw attribution weights or fuzzy inference states as in LINDA-BN and enabling straightforward expert validation.
@@
LIMREF [25] provides local, model-agnostic explanations for global TS forecasting models. It selects similar TS via DTW and generates synthetic instances via bootstrapping. Predictions guide impact rule derivation using [39]. Rules fall into four types, including counterfactual and supportive, aiding forecast adjustments. By focusing on local TS neighborhoods, LIMREF ensures rules represent global model behavior. Its focus on local fidelity complements global and progressive methods.
@@
[38] apply interpretable association rule mining to TS classification in maritime settings. They evaluate surrogate models for rule-based explanations of a neural network classifier. Each method generates decision rules, balancing fidelity and complexity. They analyze how support thresholds and discretization bins affect explanation quality. The authors highlights trade-offs between rule complexity, accuracy, and interpretability, showing the role of hyperparameter tuning in domain-specific rule optimization.
@@
[20] introduced RuleXAI [^1] https://github.com/adaa-polsl/RuleXAI, an open-source rule-based explainability library for machine learning tasks. It supports classification, regression, and survival analysis by generating elementary rule conditions — simple threshold-based splits on individual features. RuleXAI provides both global and local explanations by directly learning interpretable models from data, without relying on black-box feature attribution methods. It transforms input data into binary feature conditions, reducing preprocessing complexity and ensuring high transparency. While not designed for time series data, RuleXAI’s symbolic rule extraction aligns with decision table logic and rule-based decision systems, facilitating straightforward human interpretability.
@@
[13] explore combining numeric feature-based XAI methods in an ensemble framework, though not for TS. They propose NormEnsembleXAI, which normalizes and aggregates explanations (e.g., feature-importance, local attributions) for consistent interpretability across models. Integrating LIME, SHAP, and Integrated Gradients, they show that uniform normalization reduces attribution variability and reveals coherent explanation patterns across models.
@@
TSProto by [7] offers a complementary approach for explaining time series classifiers by converting feature attributions (e.g., SHAP or LIME) into high-level prototypes, which are then integrated into interpretable decision trees. However, it relies on a single XAI method to identify important regions and builds rules using a separate model based on SHAP-extracted features. As a result, it remains susceptible to the Rashomon effect, since in TSProto SHAP only indicates potentially relevant regions, not how the model actually uses them in decision-making.

### v4

@@ Abstract
Explaining machine learning (ML) models for time series (TS) classification remains challenging due to the difficulty of interpreting raw time series and the high dimensionality of the input space. We introduce PHARPost-hoc Attribution Rules—a unified framework that transforms numeric feature attributions from post-hoc, instance-wise explainers (e.g. LIME, SHAP) into structured, human-readable rules. These rules define human-readable intervals that indicate where and when decision-relevant segments occur and can enhance model transparency by localizing threshold-based conditions on the raw series. PHAR performs comparably to native rule-based methods, such as Anchor, while scaling more efficiently to long TS sequences and achieving broader instance coverage. A dedicated rule fusion step consolidates rule sets using strategies like weighted selection and lasso-based refinement, balancing key quality metrics: coverage, confidence, and simplicity. This fusion ensures each instance receives a concise and unambiguous rule, improving both explanation fidelity and consistency. We further introduce visualization techniques to illustrate specificity-generalization trade-offs in the derived rules. PHAR resolves conflicting and overlapping explanations—a common effect of the Rashomon phenomenon—into coherent, domain-adaptable insights. Comprehensive experiments on UCR/UEA Time Series Classification Archive demonstrate that PHAR may improve interpretability, decision transparency, and practical applicability for TS classification tasks by providing concise, human-readable rules aligned with model predictions.
@@ 1 Introduction
Explaining time series classifiers presents unique challenges compared to models for images or tabular data. Time series often lack obvious semantic features. Without domain expertise, even simple attribution outputs can be difficult to interpret, which complicates the use of standard Explainable AI (XAI) tools [theissler2022explainable]. Temporal dependencies, phase shifts, and high dimensionality further worsen the explanations. When non-specialized methods are applied, results can become unstable or overly generic [gu2024explainable]. These issues reveal a key gap in current XAI research. Unlike non-temporal models built on intuitive, semantically rich features, time series classifiers demand explanations that both capture temporal dynamics and deliver meaningful, context-aware insights tailored to sequential data. In addition, Deep Learning (DL) is widely used for (Time Series (TS)) classification, particularly in medicine and industry, because of its ability to model complex temporal patterns for tasks such because disease diagnosis and predictive maintenance. Despite strong predictive performance, these models remain difficult to interpret in practice.
@@
To address the challenges of XAI for state-of-the-art models such as DL, common approaches include numeric attribution methods such as permutation importance [altmann2010permutation], SHAP [lundberg2017unifiedapproachinterpretingmodel] and LIME [ribeiro2016why], as well as gradient-based techniques such as saliency maps [simonyan2013deep] and Layerwise Relevance Propagation (LRP) [bach2015pixel], all of which quantify feature contributions and trace relevance through the model’s computations. These approaches facilitate feature ranking, sensitivity analysis, and model debugging by producing developer-oriented outputs like importance scores or heatmaps. However, studies show that such numeric explanations often highlight broad regions of input without providing clear, context-specific insights [arsenault2024surveyexplainableartificialintelligence, jin2021deep]. In sequential data, this tendency gives rise to "ambiguous rules", where multiple time steps or variables appear equally influential and cannot be easily distinguished. Consequently, these explanations are difficult for non-experts to interpret [jeyakumar2020empirical, rojat2021explainable], and the challenge is amplified in TS classifiers, where temporal dependencies and the high dimensionality of multivariate sequences further obscure meaningful patterns [schlegel2019towards, theissler2022explainable].
@@
Anchor rule-based explanations have demonstrated improved interpretability by extracting succinct "IF–THEN" conditions that align closely with model behavior [Ribeiro_Singh_Guestrin_2018]. However, such methods are seldom applied to TS data. Furthermore, rule-based methods become computationally prohibitive as the sequence length increases [arshad2024investor]. An exhaustive search for stable "IF–THEN" conditions could be computationally infeasible, leading to a combinatorial explosion of candidate rules and temporal instability. Beyond computational infeasibility, Cognitive Load Theory, summarized by [Skulmowski2022], emphasizes reducing extraneous demands on human working memory. Analogously, simpler and unambiguous decision rules referencing a limited number of time steps minimize cognitive effort for experts, supporting interpretability and visualization by exposing time-localized, threshold-based conditions that can be inspected on the original signal.
Our motivation stems from three persistent gaps in explaining deep time series classifiers. First, both numeric attributions and native rule-based methods remain difficult to interpret on high-dimensional time series, especially when many time steps and channels contribute simultaneously. Second, most studies assess each explainer in isolation, without fusing their outputs or systematically comparing rule quality in terms of coverage, confidence, and simplicity. Third, existing work rarely overlays explanations directly on time series plots, missing an opportunity to show where and when decision-relevant segments occur.
PHAR addresses these gaps by transforming instance-level numeric attributions into compact, human-readable rules for time series classification. Rather than relying on global XAI approaches based on clustering or surrogate models [mekonnen_2024_parameterized_event_primitives], we build on local post-hoc attribution methods such as SHAP [lundberg2017unifiedapproachinterpretingmodel] and LIME [ribeiro2016why]. PHAR preserves their main strengths – precise feature ranking, post-hoc applicability, and local fidelity—while adapting them to sequential data through explicit temporal conditions. Using semi-factual reasoning, we identify minimal temporal segments in which perturbing feature values does not change the model prediction and turn these stable segments into “IF–THEN” rules. We adopt Anchor [Ribeiro_Singh_Guestrin_2018] as a baseline based on native rules and systematically compare it to rules derived from numeric attributions.
Beyond single explainers, PHAR includes a dedicated fusion step that consolidates rules from LIME, SHAP, Anchor, and potentially expert-defined rules. Fusion methods use coverage, confidence, and rule length to select a single optimized rule per instance, reducing ambiguity caused by overlapping or conflicting explanations. Finally, PHAR visualizes the resulting rules as value intervals overlaid on the original time series, allowing domain experts to inspect temporal decision boundaries directly on the signal. Although we focus on time series, the same numeric-to-rule transformation and fusion principles can be extended to other structured data, such as tabular datasets.
@@
PHAR aims to improve trust and transparency in TS classification with DL through three key contributions: (1) a pipeline that converts numeric post-hoc attributions into expert-level rule representations for TS; (2) a fusion mechanism that unifies rules from different explainers for better coherence and quality; (3) visualization of rules, as outlined in Figure 1. Each input time series is represented as a $T\times C$ matrix—$T$ time steps $t_{1},\dots,t_{T}$ by $C$ channels $c_{1},\dots,c_{C}$—and is processed by a DL classifier. Post-hoc explainers such as SHAP and LIME compute numerical attributions. In the first stage, these outputs are converted into interval rules for each instance. Rule-based explanations, such as Anchor , can be used directly. In the second stage, rule sets derived from different explainers are fused using methods such as union, weighted selection, and lasso-based refinement to yield a single coherent rule per instance. Finally, each rule is visualized as vertical interval markers overlaid on the original time-series plot, allowing domain experts to inspect and validate the model’s decision logic. Crucially, the visualization conveys information beyond the textual rule: it shows rule–signal alignment in time, overlap across channels and explainers, which supports faster error checking and expert validation.
The remainder of this paper is structured as follows: Section 2 reviews related work and explains our motivation. Section 3 details our framework, covering the transformation from numeric-to-rule (Section 3.3) and the fusion explanations (Section 3.4). Section 4 evaluates coverage, confidence, and other metrics. Section 5 summarizes findings and future directions. The Appendix A provides supplementary tables, figures, and extended results, while "Supplementary Materials" (Section B) contains statical results for Critical Difference Diagrams.
@@ 2 Related work
In this section, we review the state of the art in time series explainability that informs our PHAR framework. Existing methods follow two main lines. First, surrogate models, such as decision trees, approximate black-box predictors, and induce rule-based intervals over time. Second, post-hoc attribution methods such as LIME [ribeiro2016why], SHAP [lundberg2017unifiedapproachinterpretingmodel], and saliency maps provide numeric importance scores but often fail to enforce temporal contiguity or capture long-range dependencies in time series. Dynamic time warping (DTW) is a classical similarity measure that aligns sequences of different lengths or phases. DTW-based clustering has been used to identify prototypical patterns, offering structural insight into recurrent temporal shapes. In parallel, fuzzy rules, decision tree splits, and statistical impact rules have been proposed to improve the interpretability of sequential models, typically trading off the complexity of the rules, local fidelity, and temporal coherence. However, most of these approaches study each explainer in isolation and emphasize either local or global views, with limited support for integrating multiple explainers into a coherent description of model behavior. In what follows, we first examine rule-based explanation methods tailored to sequential data and then survey strategies for fusing outputs from multiple explainers.
@@
Explaining TS classifiers is challenging, especially for complex models.  [mekonnen_2024_parameterized_event_primitives] proposed a post-hoc method using Parameterized Event Primitives (PEPs) to explain TS models. Their approach detects key events (trends, extrema), clusters them, and derives rules via decision trees. These rules capture sequential patterns, supporting fidelity to model predictions and human interpretability, while acknowledging that such rules are proxies of model behavior rather than causal explanations. This method extracts meaningful patterns while maintaining temporal dependencies and lays the groundwork for hybrid explanations that address temporal dependencies.
@@
[SpinnatoFrancesco2023UATS] propose a TS explanation framework combining saliency maps, example-based examples, and rule-based methods. Saliency maps highlight key regions, instance-based methods provide prototypes and counterfactuals, and rule-based methods define logical conditions on subsequences. Integrating these methods bridges local and global explanations. This hybrid approach can improve the interpretability and temporal coherence of the resulting explanations by constraining rules to compact time intervals and simple inequalities on the raw series.
@@
Decision rules improve interpretability in TS forecasting, especially in finance.  [arsenault2024surveyexplainableartificialintelligence] review decision trees, fuzzy trees, and Mamdani-type rules.  [silva2021c45fuzzy] introduced FDT-FTS, which combines fuzzy logic and decision trees for multivariate TS forecasting. Mamdani-type rules use fuzzy logic and linguistic variables for human-readable explanations.  [xie2021neuralfuzzy] integrated them into the Neural Fuzzy Hammerstein-Wiener network. These methods balance predictive accuracy with human-readable simplicity of rules (few conditions, short intervals), supporting TS analysis. Next, we discuss the decision rules for progressive decision-making.
@@
The Belief–Rule–Based (BRB) framework [NimmySoniaFarhana2023] overcomes key weaknesses of post-hoc methods such as SHAP [lundberg2017unifiedapproachinterpretingmodel], LIME [ribeiro2016why], and LINDA-BN [Moreira2021LINDA] by structuring explanations into human-readable rules. In the Knowledge Graph Module (KGM), expert input is formalized into a causal graph that captures both temporal order and inter-feature dependencies, addressing SHAP’s and LIME’s tendency to treat each timestep or feature independently and thus lose important sequence information. The Knowledge Propagation Module (KPM) then propagates belief distributions through this graph to extrapolate feature trajectories, inherently managing missing data and uncertainty without relying on expensive local sampling that can render SHAP and LIME at times unstable. Finally, the Feature Evaluation Module (FEM) synthesizes these propagated beliefs into concise "IF–THEN" rules that highlight the most influential intervals, removing the need to interpret raw attribution weights or fuzzy inference states as in LINDA-BN and enabling straightforward expert validation.
@@
LIMREF [RajapakshaDilini2022LLIM] provides local model-agnostic explanations for global TS forecasting models by selecting similar TS via DTW and generating synthetic instances via bootstrapping. Predictions guide the derivation of the impact rule using [Webb2001]. Rules fall into four types, including counterfactual and supportive, which aid in forecast adjustments. By focusing on local TS neighborhoods, LIMREF ensures that rules represent global model behavior. Its focus on local fidelity complements global and progressive methods.
@@
[Veerappa_2021_Maritime] apply the interpretation of the association rule to the classification of TS in maritime settings. They evaluate surrogate models for rule-based explanations of a neural network classifier. Each method generates decision rules, balancing fidelity and complexity. They analyze how support thresholds and discretization bins affect explanation quality. The authors highlight trade-offs between rule complexity, accuracy, and interpretability, showing the role of hyperparameter tuning in domain-specific rule optimization.
@@
[RuleXAI_MACHA2022101209] introduced RuleXAI [^1] https://github.com/adaa-polsl/RuleXAI, an open-source rule-based explainability library for machine learning tasks. It supports classification, regression, and survival analysis by generating elementary rule conditions — simple threshold-based splits on individual features. RuleXAI provides both global and local explanations by directly learning interpretable models from data, without relying on black-box feature attribution methods. It transforms input data into binary feature conditions, reducing preprocessing complexity and ensuring high transparency. Although not designed for time series data, RuleXAI’s symbolic rule extraction aligns with decision table logic and rule-based decision systems, facilitating straightforward human interpretability.
@@
[hryniewskaguzik2024normensemblexaiunveilingstrengthsweaknesses] explore the combination of XAI methods based on numeric characteristics in an ensemble framework, although not for TS. They propose NormEnsembleXAI, which normalizes and aggregates explanations (e.g. feature-importance, local attributions) for consistent interpretability across models. Integrating LIME, SHAP, and Integrated Gradients, they show that uniform normalization reduces attribution variability and reveals coherent explanation patterns across models.
@@
TSProto by [sbk2025tsproto] offers a complementary approach to explaining time series classifiers by converting feature attributions (e.g. SHAP or LIME) into high-level prototypes, which are then integrated into interpretable decision trees. However, it relies on a single XAI method to identify important regions and builds rules using a separate model based on SHAP-extracted features. As a result, it remains susceptible to the Rashomon effect, since in TSProto SHAP only indicates potentially relevant regions, not how the model actually uses them in decision-making.